OpenStudy (anonymous):

A=\frac{1}{2}bh for b

OpenStudy (anonymous):

\[A=\frac{ 1 }{ 2 }bh\]

OpenStudy (anonymous):

a = 1/2bh Solve for b. To solve for a variable means to isolate the variable on one side of the equation. NOTE: We can write (1/2)(bh) as one fraction like this (bh/2). They mean the same thing. We now have this: a = (bh/2) We multiply both sides of the fractional equation by 2. This is done to remove the fraction. Doing this we now have a linear equation: 2a = bh We now divide both sides of the equation by h to find the value of b. Final answer: (2a/h) = b Did you follow?

OpenStudy (anonymous):

yes, I followed. thank you

OpenStudy (anonymous):

ok

OpenStudy (jiteshmeghwal9):

\[\LARGE{A=\frac{1}{2}bh}\]take 1/2 @ leftside then it will become\[\LARGE{\frac{1}{2}A=bh}\]now similarly take 'h' @ left side